OpenStudy (anonymous):

Find the derivative of f(x)=sqrt((1)/(x^9))

OpenStudy (unklerhaukus):

\[f(x)=\sqrt{\frac{1}{x^9}}=x^{-9/2}\]

OpenStudy (unklerhaukus):

\[f'(x)=\frac{-9}{2}x^{-11/2}\]

OpenStudy (anonymous):

Thanks!